# TWISTED CONJUGACY IN DIHEDRAL ARTIN GROUPS I: TORUS KNOT GROUPS

### v3

@@
Abstract
In this paper we provide an alternative solution to a result by Juhász [14], that the twisted conjugacy problem for odd dihedral Artin groups is solvable, that is, groups with presentation $G(m)=\langle a,b\mid{}_{m}(a,b)={}_{m}(b,a)\rangle$, where $m\geq 3$ is odd, and (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$, is solvable. Our solution provides an implementable linear time algorithm, by considering an alternative group presentation to that of a torus knot group, and working with geodesic normal forms. An application of this result is that the conjugacy problem is solvable in extensions of odd dihedral Artin groups.
2020 Mathematics Subject Classification: 20F10, 20F36.
[…]
@@ INTRODUCTION
As an application of Section 1, we consider the conjugacy problem in extensions of odd dihedral Artin groups, using a criteria from [3] (see 5.1). This criteria requires understanding orbit decidability for subgroups $A\leq\mathrm{Aut}(G)$, for some group $G$. The orbit decidability problem asks whether we can determine if for two elements $u,v\in G$, there exists an automorphism $\phi\in A$ such that $v$ is conjugate to $\phi(u)$. To apply the criteria from [3], we require what is known as the ‘action subgroup’ of odd dihedral Artin groups to be orbit decidable. For odd dihedral Artin groups, we prove a stronger statement, by showing that all subgroups of the automorphism group of odd dihedral Artin groups are orbit decidable.
@@
thmorbit Every finitely generated subgroup $A\leq\mathrm{Aut}(G(m))$, when $m$ is odd, is orbit decidable. Combined with Section 1, we can find new examples of groups with solvable conjugacy problem.
@@
thmextension Let $G=G(m)\rtimes H$ be an extension of an odd dihedral Artin group by a finitely generated group $H$ which satisfies conditions $(ii)$ and $(iii)$ from 5.1 (e.g. let $H$ be torsion-free hyperbolic). Then $G$ has decidable conjugacy problem. We note that the conjugacy problem is solvable for virtual dihedral Artin groups [8, Corollary 3.2], which proves Section 1 in the case where $H$ is finite.
@@
Decision problems and twisted conjugacy
Let $X$ be a finite set, and let $X^{\ast}$ be the set of all finite words over $X$. For a group $G$ generated by $X$, we use $u=v$ to denote equality of words in $X^{\ast}$, and $u=_{G}v$ to denote equality of the group elements represented by $u$ and $v$. We let $l(w)$ denote the word length of $w$ over $X$. For a group element $g\in G$, we define the length of $g$, denoted $|g|_{X}$, to be the length of a shortest representative word for the element $g$ over $X$. A word $w\in X^{\ast}$ is geodesic if $l(w)=|\pi(w)|_{X}$, where $\pi\colon X^{\ast}\rightarrow G$ is the natural projection. If there exists a unique word $w$ of minimal length representing $g$, then we say $w$ is a unique geodesic. Otherwise, $w$ is a non-unique geodesic.
Definition 2.1.
@@
Remark 2.3.
Let $\psi\circ\phi\in\mathrm{Aut}(G)$, where $\psi\in\mathrm{Inn}(G)$, $\phi\in\mathrm{Out}(G)$ are fixed. If $\phi$ is trivial, then the $\mathrm{TCP}_{\psi}(G)$ is equivalent to solving the $\mathrm{CP}(G)$ by the following relations:
@@
$v=_{G}(\psi\circ\phi)(w)^{-1}uw\Leftrightarrow v=_{G}g^{-1}\phi(w)^{-1}guw\Leftrightarrow gv=_{G}\phi(w)^{-1}(gu)w,$
for some fixed $g\in G$. Therefore if $\phi\in\mathrm{Out}(G)$ is non-trivial, then the $\mathrm{TCP}_{\psi\circ\phi}(G)$ is equivalent to solving the $\mathrm{TCP}_{\phi}(G)$, for some fixed $\phi\in\mathrm{Out}(G)$.
Definition 2.4.
@@
Odd dihedral Artin groups
Definition 2.5.
@@
where (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$.
We first note the following result, which solves the $\mathrm{TCP}_{\phi}(G(m))$ in the case where the automorphism $\phi\in\mathrm{Aut}(G(m))$ is trivial.
Theorem 2.6.
@@
Definition 2.7.
Let $G(m)$ denote the dihedral Artin group, with presentation given in Eq. 1. For $m=2k+1$ for some $k\in\mathbb{Z}_{\geq 1}$,
$G(m)\cong\langle x,y\mid x^{2}=y^{m}\rangle,$ (2)
by setting $x={}_{m}(a,b),y=ab$. Here $G(m)$ is isomorphic to a torus knot group. For notation we let $X=\{x,y\}^{\pm}$, which we refer to as the free product generating set.
@@
In particular, $\mathrm{Aut}(G(m))=\mathrm{Inn}(G(m))\sqcup\mathrm{Inn}(G(m))\cdot\phi$ where $\phi\in\mathrm{Out}(G(m))$ inverts all generators, i.e.
@@
For the remainder of this section, we assume $\phi\in\mathrm{Out}(G(m))$ is as in Eq. 3. We let $\mathsf{rev}(w)$ denote the word $w\in X^{\ast}$ written in reverse, and note that for any word $w\in X^{\ast}$, $\phi(w)^{-1}=\mathsf{rev}(w)$. Indeed if $w=w^{a_{1}}_{i_{1}}\dots w^{a_{n}}_{i_{n}}\in X^{\ast}$, where each $w_{i_{j}}\in X$ and $a_{j}\in\mathbb{Z}_{\neq 0}$ $(1\leq j\leq n)$, then $\phi(w)=w^{-a_{1}}_{i_{1}}\dots w^{-a_{n}}_{i_{n}}$, and so $\phi(w)^{-1}=w^{a_{n}}_{i_{n}}\dots w^{a_{1}}_{i_{1}}=\mathsf{rev}(w)$. Therefore checking if two elements are twisted conjugate reduces to checking if $v=_{G}\mathsf{rev}(w)uw$.
@@
Definition 2.10.
Let $u=x^{a_{1}}y^{b_{1}}\dots y^{b_{\tau}}\Delta^{c}\in X^{\ast}$ be geodesic. We say $u$ is Garside-free if $c=0$.
Let $u\in X^{\ast}$ be a geodesic, where $l(u)>1$. We can write $u$ in the form $u=u_{1}u_{2}$ for some non-empty geodesic words $u_{1},u_{2}\in X^{\ast}$. We say $u_{1}$ is a proper prefix of $u$, and $u_{2}$ is a proper suffix of $u$. We let $\mathcal{P}(u)$ and $\mathcal{S}(u)$ be the set of all possible proper prefixes and suffixes of $u$ respectively. With these definitions, we can define an equivalent notion of cyclic permutations, with respect to twisted conjugacy.
Definition 2.11.
Let $u\in X^{\ast}$ be a geodesic, where $l(u)>1$. Let $u=u_{1}u_{2}$ for some $u_{1}\in\mathcal{P}(u),\;u_{2}\in\mathcal{S}(u)$. Let $\phi\in\mathrm{Out}(G(m))$ as defined in Eq. 3. We define a $\phi$-cyclic permutation of a prefix of $u$ to be the following operation on $u$:
@@
ALGORITHM FOR OUTER AUTOMORPHISMS
Our main goal is to solve the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$, where $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3. We will then generalise this algorithm, using 2.3, to solve the twisted conjugacy problem $\mathrm{TCP}(G(m))$.
We will first show that when $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3, then every geodesic is twisted conjugate to a minimal length twisted conjugacy representative which is Garside free. Moreover, for two minimal representatives which are twisted conjugate, we will show that they are related by a finite sequence of $\phi$-cyclic permutations and equivalent geodesics. Since $\phi$ is length-preserving, this will lead to our basic algorithm, which we later improve to give a linear time algorithm.
Garside free representatives
Definition 3.1.
Let $u\in X^{\ast}$ be geodesic. We say $u$ is of Type (3) if $u$ is not of Type (1) or (2) (see Table 2). In particular, Type (3) geodesics are Garside free (recall 2.10). We let $\overline{(3)}\subset(3)$ denote the set of all words $u\in(3)$, such that $u$ starts and ends with opposite letters.
Our first step is to prove the following result, which shows that any geodesic $u\in X^{\ast}$ can be transformed to a geodesic $u^{\prime}\in X^{\ast}$, which is Garside free, such that $u\sim_{\phi}u^{\prime}$.
Proposition 3.2.
@@
Now suppose $v=y^{t+mc}=_{G.}y^{t}\Delta^{c}$. If $c=2d$, for some $d\in\mathbb{Z}$, then $v=_{G}y^{t}\Delta^{2d}\xleftrightarrow{\phi}\Delta^{-d}y^{t}\Delta^{d}=_{G}y^{t}.$ If $t$ is even, then $y^{t}\xleftrightarrow{\phi}y^{-t/2}y^{t/2}=_{G}\varepsilon$. Similarly if $t$ is odd, then $y^{t}\xleftrightarrow{\phi}y^{\pm 1}$. Otherwise, if $c=2d+1$, then $v=y^{t}\Delta^{2d+1}\xleftrightarrow{\phi}\Delta^{-d}y^{t}\Delta\Delta^{d}=_{G}y^{t}\Delta=y^{t+m}.$ If $t$ is odd, then $t+m$ is even, and so $y^{t+m}\xleftrightarrow{\phi}\varepsilon$. Similarly if $t$ is even, then $t+m$ is odd and so $y^{t+m}\xleftrightarrow{\phi}y^{\pm 1}$. ∎
To prove 3.2, we will construct a set of rewriting rules which can be repeatedly applied to a geodesic $u\in X^{\ast}$ to obtain an element $u^{\prime}\in\overline{(3)}$. These rules are invariant under $\sim_{\phi}$, that is, after applying a sequence of these rules to $u$, we obtain a word $u^{\prime}\in\overline{(3)}$ such that $u\sim_{\phi}u^{\prime}$. Let $v\in X^{\ast}$ such that $v$ is Garside free, let $x,y\in X$ and $\epsilon,\epsilon_{1},\epsilon_{2}\in\{\pm 1\}$. We define our set of rewrite rules as follows.
@@
Proof of 3.2.
Assume $u\not\in\overline{(3)}$. If $u$ contains only $x$ or $y$ factors, then we obtain $u^{\prime}\in\{\varepsilon,x^{\pm 1},y^{\pm 1}\}$, such that $u\sim_{\phi}u^{\prime}$, based on the conditions given in Lemma 3.3.
@@
  1. 1.
    If $c\neq 0$, apply (R1) to obtain an element of the form $v$ or $v\Delta$.
  2. 2.
    If $v$ starts and ends with the same letter, then repeatedly apply (R2)-(R3), to obtain an element of the form $v$ or $v\Delta^{\pm 1}$, where $v\in\overline{(3)}$. If, after any application of (R2)-(R3), we obtain an element of the form $v\Delta^{\pm 2}$, then return to Step 1 and apply (R1).
  3. 3.
    If, after Step 2, we obtain $v\Delta^{\pm 1}$ where $v\in\overline{(3)}$, then apply (R4) or (R5) to obtain an element $v\in\overline{(3)}$.
By 3.4, $u$ is twisted conjugate to the output of this sequence of rewrite rules. Since (R1)-(R5) either preserves or decreases word length, then $l(u^{\prime})\leq l(u)$. ∎
Twisted cyclic geodesics
We now consider which elements of Type $\overline{(3)}$, as defined in 3.1, are minimal length up to twisted cyclic permutations.
Definition 3.5.
Let $G=G(m)$ be a group with presentation as in Eq. 2. We define the set of twisted cyclic geodesics, denoted $\mathsf{CycGeo}_{\phi}(G,X)$, to be the set of all geodesics $u\in X^{\ast}$, such that every $\phi$-cyclic permutation of $u$ is also geodesic.
Let $\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)=\mathsf{CycGeo}_{\phi}(G,X)\cap\;\text{Type}\;\overline{(3)}$ be the set of all twisted cyclic geodesics of Type $\overline{(3)}$. The following example highlights that when we apply a $\phi$-cyclic permutation to $u\in\overline{(3)}$, we can obtain a word which is no longer geodesic.
Example 3.6.
Suppose $m=5$, and consider $u=y^{-3}x^{-1}\in X^{\ast}$, which is geodesic of length 4. We can take a twisted cyclic permutation of $u$ by applying the $\phi$-cyclic permutation $u=y^{-3}x^{-1}\xleftrightarrow{\phi}\phi\left(x^{-1}\right)y^{-3}=xy^{-3}=u^{\prime}.$ Here $u^{\prime}$ is not geodesic, since $u^{\prime}=_{G}x^{-1}y^{2}$, which is of length 3.
To consider when words of Type $\overline{(3)}$ lie in $\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)$, we first consider when geodesics of Type (3) are geodesic.
Definition 3.7.
For $u=x^{a_{1}}y^{b_{1}}\dots x^{a_{\tau}}y^{b_{\tau}}\in(3)$, define

### v5

@@
Abstract.
In this paper we provide an alternative solution to a result by (Juhász 2011), that the twisted conjugacy problem for odd dihedral Artin groups is solvable, where an odd dihedral Artin group is a group with presentation $G(m)=\langle a,b\mid{}_{m}(a,b)={}_{m}(b,a)\rangle$, where $m\geq 3$ is odd, and (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$. Our solution provides an implementable linear time algorithm, by considering an alternative group presentation to that of a torus knot group, and working with geodesic normal forms. An application of this result is that the conjugacy problem is solvable in extensions of odd dihedral Artin groups.
@@ 1. Introduction
As an application of Theorem 1, we consider the conjugacy problem in extensions of odd dihedral Artin groups, using a criteria from [3] (see Theorem 23). This criteria requires understanding orbit decidability for subgroups $A\leq\mathrm{Aut}(G)$, for some group $G$. The orbit decidability problem asks whether we can determine if for two elements $u,v\in G$, there exists an automorphism $\phi\in A$ such that $v$ is conjugate to $\phi(u)$. To apply the criteria from [3], we require what is known as the ‘action subgroup’ of odd dihedral Artin groups to be orbit decidable. For odd dihedral Artin groups, we prove a stronger statement, by showing that all subgroups of the automorphism group of odd dihedral Artin groups are orbit decidable.
@@
thmorbit Every finitely generated subgroup $A\leq\mathrm{Aut}(G(m))$, when $m$ is odd, is orbit decidable. Combined with Theorem 1, we can find new examples of groups with solvable conjugacy problem.
@@
thmextension Let $G=G(m)\rtimes H$ be an extension of an odd dihedral Artin group by a finitely generated group $H$ which satisfies conditions $(ii)$ and $(iii)$ from Theorem 23 (e.g. let $H$ be torsion-free hyperbolic). Then $G$ has decidable conjugacy problem. We note that the conjugacy problem is solvable for virtual dihedral Artin groups [8, Corollary 3.2], which proves Theorem 1 in the case where $H$ is finite.
@@ 2. Preliminaries
2.1. Decision problems and twisted conjugacy
Let $X$ be a finite set, and let $X^{\ast}$ be the set of all finite words over $X$. For a group $G$ generated by $X$, we use $u=v$ to denote equality of words in $X^{\ast}$, and $u=_{G}v$ to denote equality of the group elements represented by $u$ and $v$. We let $l(w)$ denote the word length of $w$ over $X$. For a group element $g\in G$, we define the length of $g$, denoted $|g|_{X}$, to be the length of a shortest representative word for the element $g$ over $X$. A word $w\in X^{\ast}$ is geodesic if $l(w)=|\pi(w)|_{X}$, where $\pi\colon X^{\ast}\rightarrow G$ is the natural projection. If there exists a unique word $w$ of minimal length representing $g$, then we say $w$ is a unique geodesic. Otherwise, $w$ is a non-unique geodesic. We let $\varepsilon$ denote the empty word over $X$ representing the identity element of $G$.
{defi}
@@ 2.1. Decision problems and twisted conjugacy
Remark 1.
Let $\psi\in\mathrm{Aut}(G)$, and let $[\psi]$ denote the equivalence class of $\psi$ in $\mathrm{Out}(G)$. We can write $\psi$ in the form $\psi=\iota_{g}\phi$, where $\iota_{g}\in\mathrm{Inn}(G)$ denotes conjugation by $g$, and $\phi\in[\psi]$. If $[\psi]$ is trivial, then the $\mathrm{TCP}_{\psi}(G)$ is equivalent to solving the $\mathrm{CP}(G)$ by the following relations:
@@
$v=_{G}(\iota_{g}\phi)(w)^{-1}uw\Leftrightarrow v=_{G}g^{-1}\phi(w)^{-1}guw\Leftrightarrow gv=_{G}\phi(w)^{-1}(gu)w,$
for some fixed $g\in G$. In particular, if $[\psi]$ is non-trivial, then the $\mathrm{TCP}_{\psi}(G)$ is equivalent to solving the $\mathrm{TCP}_{\phi}(G)$ for some fixed $\phi\in[\psi]$.
{defi}
@@
2.2. Odd dihedral Artin groups
{defi}
@@ 2.2. Odd dihedral Artin groups
where (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$. We first note the following result, which solves the $\mathrm{TCP}_{\phi}(G(m))$ in the case where the automorphism $\phi\in\mathrm{Aut}(G(m))$ is trivial.
Theorem 2.
@@
{defi}
Let $m=2k+1$, where $k\in\mathbb{Z}_{\geq 1}$. We define
$P=\langle x,y\mid x^{2}=y^{m}\rangle$ (2)
to be the presentation of a torus-knot group. For notation we let $X=\{x,y\}^{\pm}$, which we refer to as the free product generating set.
Lemma 3.
When $m\geq 3$ is odd, the presentations given in Eq. 1 and Eq. 2 represent isomorphic groups.
Proof 2.1.
We apply a sequence of Tietze transformations to $G(m)$ to obtain $P$ as follows. First, set $x={}_{m}(a,b)$ and $y=ab$, which gives us
$G(m)\cong\langle a,b,x,y\mid{}_{m}(a,b)={}_{m}(b,a),x={}_{m}(a,b),y=ab\rangle.$
Since $x=by^{k}=y^{k}a$, then $x^{2}=y^{k}aby^{k}=y^{2k+1}$, and so our relations can be rewritten as
$G(m)\cong\langle a,b,x,y\mid x={}_{m}(a,b)={}_{m}(b,a),y=ab,x^{2}=y^{2k+1}\rangle.$
Next we remove the generator $a$, by rewriting our relation $y=ab$ as $a=yb^{-1}$. Removing redundant relations gives us
$G(m)\cong\langle b,x,y\mid x=y^{k+1}b^{-1}=by^{k},x^{2}=y^{2k+1}\rangle.$
Finally we remove the generator $b$ in a similar way by rewriting the relation $x=by^{k}$, and removing redundant relations. This leaves us with $G(m)\cong P$ as required.
@@
In particular, $\mathrm{Aut}(G(m))=\mathrm{Inn}(G(m))\sqcup\mathrm{Inn}(G(m))\cdot\phi$ where $\phi$ inverts all generators, i.e.
@@
For the remainder of this section, we assume $\phi\in\mathrm{Aut}(G(m))$ as in Eq. 3. We let $\mathsf{rev}(w)$ denote the word $w\in X^{\ast}$ written in reverse, and note that for any word $w\in X^{\ast}$, $\phi(w)^{-1}=\mathsf{rev}(w)$. Indeed if $w=w^{a_{1}}_{i_{1}}\dots w^{a_{n}}_{i_{n}}\in X^{\ast}$, where each $w_{i_{j}}\in X$ and $a_{j}\in\mathbb{Z}_{\neq 0}$ $(1\leq j\leq n)$, then $\phi(w)=w^{-a_{1}}_{i_{1}}\dots w^{-a_{n}}_{i_{n}}$, and so $\phi(w)^{-1}=w^{a_{n}}_{i_{n}}\dots w^{a_{1}}_{i_{1}}=\mathsf{rev}(w)$. Therefore checking if two elements are twisted conjugate reduces to checking if $v=_{G}\mathsf{rev}(w)uw$.
@@
{defi}
Let $u=x^{a_{1}}y^{b_{1}}\dots y^{b_{\tau}}\Delta^{c}\in X^{\ast}$ be geodesic. We say $u$ is Garside-free if $c=0$. Let $u\in X^{\ast}$ be a geodesic, where $l(u)>1$. We can write $u$ in the form $u=u_{1}u_{2}$ for some non-empty geodesic words $u_{1},u_{2}\in X^{\ast}$. We say $u_{1}$ is a proper prefix of $u$, and $u_{2}$ is a proper suffix of $u$. We let $\mathcal{P}(u)$ and $\mathcal{S}(u)$ be the set of all possible proper prefixes and suffixes of $u$ respectively. With these definitions, we can define an equivalent notion of cyclic permutations, with respect to twisted conjugacy. {defi} Let $u\in X^{\ast}$ be a geodesic, where $l(u)>1$. Let $u=u_{1}u_{2}$ for some $u_{1}\in\mathcal{P}(u),\;u_{2}\in\mathcal{S}(u)$. Let $\phi\in\mathrm{Aut}(G(m))$ as defined in Eq. 3. We define a $\phi$-cyclic permutation of a prefix of $u$ to be the following operation on $u$:
@@
3. Algorithm for Outer Automorphisms
Our main goal is to solve the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$, where $\phi\in\mathrm{Aut}(G(m))$ as in Eq. 3. We will then generalise this algorithm, using Remark 1, to solve the twisted conjugacy problem $\mathrm{TCP}(G(m))$.
We will first show that when $\phi\in\mathrm{Aut}(G(m))$ as in Eq. 3, then every geodesic is twisted conjugate to a minimal length twisted conjugacy representative which is Garside free. Moreover, for two minimal representatives which are twisted conjugate, we will show that they are related by a finite sequence of $\phi$-cyclic permutations and equivalent geodesics. Since $\phi$ is length-preserving, this will lead to our basic algorithm, which we later improve to give a linear time algorithm.
3.1. Garside free representatives
{defi}
Let $u\in X^{\ast}$ be geodesic. We say $u$ is of Type (3) if $u$ is not of Type (1) or (2) (see Table 2). In particular, Type (3) geodesics are Garside free (recall Definition 2.2). We let $\overline{(3)}\subset(3)$ denote the set of all words $u\in(3)$, such that $u$ starts and ends with opposite letters. Our first step is to prove the following result, which shows that any geodesic $u\in X^{\ast}$ can be transformed to a geodesic $u^{\prime}\in X^{\ast}$, which is Garside free, such that $u\sim_{\phi}u^{\prime}$.
Proposition 6.
@@ 3.1. Garside free representatives
Now suppose $v=y^{t+mc}=_{G.}y^{t}\Delta^{c}$. If $c=2d$, for some $d\in\mathbb{Z}$, then $v=_{G}y^{t}\Delta^{2d}\xleftrightarrow{\phi}\Delta^{-d}y^{t}\Delta^{d}=_{G}y^{t}.$ If $t$ is even, then $y^{t}\xleftrightarrow{\phi}y^{-t/2}y^{t/2}=_{G}\varepsilon$. Similarly if $t$ is odd, then $y^{t}\xleftrightarrow{\phi}y^{\pm 1}$. Otherwise, if $c=2d+1$, then $v=y^{t}\Delta^{2d+1}\xleftrightarrow{\phi}\Delta^{-d}y^{t}\Delta\Delta^{d}=_{G}y^{t}\Delta=y^{t+m}.$ If $t$ is odd, then $t+m$ is even, and so $y^{t+m}\xleftrightarrow{\phi}\varepsilon$. Similarly if $t$ is even, then $t+m$ is odd and so $y^{t+m}\xleftrightarrow{\phi}y^{\pm 1}$.
To prove Proposition 6, we will construct a set of rewriting rules which can be repeatedly applied to a geodesic $u\in X^{\ast}$ to obtain an element $u^{\prime}\in\overline{(3)}$. These rules are invariant under $\sim_{\phi}$, that is, after applying a sequence of these rules to $u$, we obtain a word $u^{\prime}\in\overline{(3)}$ such that $u\sim_{\phi}u^{\prime}$. Let $v\in X^{\ast}$ such that $v$ is Garside free, let $x,y\in X$ and $\epsilon,\epsilon_{1},\epsilon_{2}\in\{\pm 1\}$. We define our set of rewrite rules as follows.
@@
Proof 3.3 (Proof of Proposition 6).
Assume $u\not\in\overline{(3)}$. If $u$ contains only $x$ or $y$ factors, then we obtain $u^{\prime}\in\{\varepsilon,x^{\pm 1},y^{\pm 1}\}$, such that $u\sim_{\phi}u^{\prime}$, based on the conditions given in Lemma 7.
@@
  1. (1)
    If $c\neq 0$, apply (R1) to obtain an element of the form $u_{1}$ or $u_{1}\Delta$.
  2. (2)
    If $u_{1}$ starts and ends with the same letter, then repeatedly apply (R2)-(R3), to obtain an element of the form $u_{1}$ or $u_{1}\Delta^{\pm 1}$, where $u_{1}\in\overline{(3)}$. If, after any application of (R2)-(R3), we obtain an element of the form $u_{1}\Delta^{\pm 2}$, then return to Step 1 and apply (R1).
  3. (3)
    If, after Step 2, we obtain $u_{1}\Delta^{\pm 1}$ where $u_{1}\in\overline{(3)}$, then apply (R4) or (R5) to obtain an element $u_{1}\in\overline{(3)}$.
By Proposition 8, $u$ is twisted conjugate to the output $u^{\prime}$ of this sequence of rewrite rules. Since (R1)-(R5) either preserves or decreases word length, then $l(u^{\prime})\leq l(u)$.
3.2. Twisted cyclic geodesics
We now consider which elements of Type $\overline{(3)}$, as defined in Definition 3.1, are minimal length up to twisted cyclic permutations.
{defi}
Let $G=G(m)$ be a group with presentation as in Eq. 2. We define the set of twisted cyclic geodesics, denoted $\mathsf{CycGeo}_{\phi}(G,X)$, to be the set of all geodesics $u\in X^{\ast}$, such that every $\phi$-cyclic permutation of $u$ is also geodesic. Let $\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)=\mathsf{CycGeo}_{\phi}(G,X)\cap\;\text{Type}\;\overline{(3)}$ be the set of all twisted cyclic geodesics of Type $\overline{(3)}$. The following example highlights that when we apply a $\phi$-cyclic permutation to $u\in\overline{(3)}$, we can obtain a word which is no longer geodesic. {exa} Suppose $m=5$, and consider $u=y^{-3}x^{-1}\in X^{\ast}$, which is geodesic of length 4. We can take a twisted cyclic permutation of $u$ by applying the $\phi$-cyclic permutation $u=y^{-3}x^{-1}\xleftrightarrow{\phi}\phi\left(x^{-1}\right)y^{-3}=xy^{-3}=u^{\prime}.$ Here $u^{\prime}$ is not geodesic, since $u^{\prime}=_{G}x^{-1}y^{2}$, which is of length 3. To consider when words of Type $\overline{(3)}$ lie in $\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)$, we first consider when words of Type (3) are geodesic. {defi} For $u=x^{a_{1}}y^{b_{1}}\dots x^{a_{\tau}}y^{b_{\tau}}\in(3)$, define